OpenStudy (anonymous):

A right triangle has base (x - 5) units and height (4x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle? Part B: What is the area of the triangle? Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials.

OpenStudy (campbell_st):

well for A using pythagoras' theorem \[h^2 = (4x -10)^2 + (x -5)^2\] for B the area \[A = \frac{1}{2} \times (x -5)(4x -10)\]

OpenStudy (anonymous):

I understand all that but idk what c is

OpenStudy (campbell_st):

all that is is says is that if you multiply 2 polynomials then you get a polynomial as an answer... if you add 2 polynomials you still get a polynomial. so using pythagoras' theorem squaring each polynomial will result in a degree 2 polynomial then adding the 2 degree 2 polynomials still results in a polynomial so it is a closed operation. the area question, you have degree 0 polynomial, degree 1 polynomial and degree 1 polynomial so multiplying them will result in a polynomial of degree 2 hope it makes sense